A businessman has three times as much money invested at 3% as he does at 4.5%. The income from both investments for a year is $675. How much is invested at each rate.

Answer by stanbon(75887)

A businessman has three times as much money invested at 3% as he does at 4.5%. The income from both investments for a year is $675. How much is invested at each rate.
-----------------------------
Equations:
Quantity: x = 3y
Interest:0.03x + 0.045y = 675
--------------------------------
Substitute for "x":
0.03(3y) + 0.045y = 675
---
Multiply thru by 1000:
3(30y) + 45y = 675000
135y = 675000
y = $5000.00
x = (1/3)(5000) = $1666.67
================================
Cheers,
Stan H.